OpenStudy (anonymous):

calculate, using the productrule, the derivative of x (2logx) Product rule: (f(x) (g(x))' = f'(x) g(x) + f(x) g'(x) f(x) = x f'(x) = 1 g(x) = 2log(x) g'(x) = ??? I think 2Log(x) = 1 / xln2 then fill in the productrule : 1 . 2log(x) + x . (1 / xln2) = 2log(x) + x / xln2 however. the asnwer book sais it should be (1 + ln x) / (ln2)

OpenStudy (amistre64):

g' = 2/x

OpenStudy (amistre64):

of course, depending on what math class you are int, log can mean base 10, or it can be a natural log .... if base 10, then g' = 2/(xln(10))

OpenStudy (anonymous):

I still don't get that part. It's a base 10. but why is the g' 2/x because of the 2 standing in front of the log ?

OpenStudy (amistre64):

a few things to keep in mind: constants pull out, and logs have a change of base process. \[log_b(x)=\frac{ln(x)}{ln(b)}\] 1/ln(b) is just another constant. \[2log_{10}(x)=\frac{2}{ln(10)}ln(x)\]pull out the constant and derive the ln(x)